# London WC2R 2LS, UK

### v1

@@ 3.3 Worldsheet symmetry and geometry
$\widehat{\nabla}$-covariantly constant forms on the spacetime $M^{D}$ generate symmetries in sigma model actions, like for example that of (2) HowePap . More specifically, if the sigma model target space $M^{D}$ admits a $\widehat{\nabla}$ covariantly constant $(\ell+1)$-form $L$, i.e.
^$\widehat{\nabla}_{I}L_{J_{1}\dots J_{\ell+1}}=0~{},$ (81)
@@
where $\varepsilon_{L}$ is the infinitesimal parameter whose Grassmannian parity is chosen such that the variation $\delta_{L}X$ is even. It turns out that the conditions imposed on $F$ in (57), (66), (72) and (79) all follow from the condition (83) that arises from the requirement of invariance of heterotic sigma model action under the transformations (82) generated by the fundamental forms of the holonomy groups $SU(n)$, $Sp(k)$, $G_{2}$ and $\mathrm{Spin}(7)$ of $\widehat{\nabla}$ connection. This justifies the restrictions we have put on the curvature of the gauge sector in the discussion of geometries that satisfy the scale invariance conditions.
The commutator, $[\delta_{L},\delta_{P}]X$ of two such transformations on $X$ generated by the $(\ell+1)$-form $L$ and the $(p+1)$-form $P$ on $M$ is rather involved to be described here, see HPS for a detailed exposition. Even in the $G_{2}$ and ${\mathrm{S}pin}(7)$ cases, where the symmetries are generated by the fundamental forms that exhibit special algebraic properties, the expressions are rather involved especially in the case that $H\not=0$. In the $G_{2}$ and ${\mathrm{S}pin}(7)$, the algebra closes as a W-algebra, i.e. the structure constants of the algebra depend on conserved currents. The commutator on $\Psi$ is
$[\delta_{L},\delta_{P}]\Psi_{-}^{a}=-A_{K}{}^{a}{}_{b}[\delta_{L},\delta_{P}]X^{K}-F_{IJ}{}^{a}{}_{b}\delta_{L}X^{I}\delta_{P}X^{J}\Psi_{-}^{b}~{}.$ (84)
@@
The simplest cases to consider are those for which the sigma-model target space admits either a KT structure $I$ (^$\widehat{\nabla}I=0$), or an HKT structure $I_{r}$ (^$\widehat{\nabla}I_{r}=0$), then the action (2) is invariant under (0,2) or (0,4) worldsheet supersymmetry transformations [^11] There is an extensive literature, see for example Curtrightfreedman ; LAGDF ; Curtrightzachos ; Howesierra ; JGCHMR ; Buscher ; Braaten ; Hullreview ; HoweGP1 , on the geometry of target spaces of 2-dimensional sigma models with $(p,q)$, $p,q\not=0$, supersymmetry following the early work of Zumino2 in four dimensions., respectively. The additional supersymmetry transformations Hullwitten are given by
@@
where the integrability condition of the complex structures $I_{r}$ has been used [^12] Vanishing of the Nijenhuis tensor of $I_{r}$ simplifies the right hand side of the commutator (87). together with the conditions (83) for $L=I_{r}$. The latter conditions imply that $F$ is a (1,1)-form with respect to each complex structure $I_{r}$. In the KT case, the second term in the right hand side of the commutator can be expressed as $(\varepsilon D_{+}\varepsilon^{\prime}-\varepsilon^{\prime}D_{+}\varepsilon)D_{+}X^{K}$, while in the HKT case it can be expressed as $\delta_{rs}(\varepsilon^{r}D_{+}\varepsilon^{\prime s}-\varepsilon^{\prime s}D_{+}\varepsilon^{r})D_{+}X^{K}-(\varepsilon^{r}D_{+}\varepsilon^{\prime s}+\varepsilon^{\prime s}D_{+}\varepsilon^{r})\epsilon_{rs}{}^{t}I_{t}{}^{K}{}_{L}D_{+}X^{L}$. In either case, the commutator (87) closes to spacetime translations and supersymmetry transformations as expected. The commutator of two transformations (85) on $\Psi$ closes on-shell to the standard $(0,q)$ supersymmetry algebra, for either $q=2$ or $q=4$.
@@
where the superfields $X$ and $\Psi$ satisfy the constraint (88). It can be shown that this action is invariant under all $(0,q)$ supersymmetry transformations provided that in addition the fibre metric $h$ is (1,1) tensor with respect to the fibre complex structures $J_{r}$, i.e. $h_{a^{\prime}b^{\prime}}J_{r}{}^{a^{\prime}}{}_{a}J_{r}{}^{b^{\prime}}{}_{b}=h_{ab}$ for either $r=1$ or $r=1,2,3$.
@@ 4.1 Comparison of conditions for scale and conformal invariance
^$\widehat{\nabla}^{J}F_{JI}{}^{a}{}_{b}+\widehat{\theta}^{J}F_{JI}{}^{a}{}_{b}=0~{},$ (91)
provided that (12) holds and the connection $\widetilde{\nabla}$ and ${\mathcal{D}}$ satisfy instanton-like conditions. A comparison of these conditions for scale invariance with the conformal invariance conditions, which are the vanishing conditions for the beta functions of $G$, $B$ and $A$, in (21), reveals that these special geometries solve the latter provided that the 1-form
@@
^$\widehat{\nabla}_{K}\widehat{V}_{L}=0~{},~{}~{}~{}\widehat{V}^{K}F_{KL}{}^{a}{}_{b}=0\,,$ (93)
where $\Phi$ is the dilaton. Therefore, $\widehat{V}$ is $\widehat{\nabla}$-covariantly constant and the directions of $F$ along $\widehat{V}$ vanish. The former condition implies that
@@
Alternatively, suppose that $\widehat{V}\not=0$. As $\widehat{V}$ is non-vanishing and $\widehat{\nabla}$-covariantly constant, it is non-vanishing everywhere on $M^{D}$. Moreover its length is constant as $\widehat{\nabla}$ is metric. The integrability condition of ^$\widehat{\nabla}\widehat{V}=0$ reveals that
@@
and the holonomy of the connection $\widehat{\nabla}$ reduces further to a subgroup of the original holonomy group stated in (1) whose geometry was summarised in sections 3.1 and 3.2. The pattern of these holonomy reductions depends on the original holonomy group and they will be explored below together with some aspects of the geometry of $M^{D}$. Here, we shall proceed to show that the couplings are invariant under the diffeomorphisms generated by $\widehat{V}$.
@@
4.2 Geometry of conformally invariant backgrounds with $\widehat{V}=0$ and spacetime supersymmetry
@@
We have demonstrated that the special geometries investigated in section 3 that satisfy the conformal invariance conditions with $\widehat{V}=0$ are conformally balanced, i.e. $\widehat{\theta}=-2{\mathrm{d}}\Phi$. The conformal balance condition has arisen before in the context of (spacetime) supersymmetric heterotic backgrounds. Here, we shall address the question whether the invariance conditions (81) and (83) on the sigma model couplings we have found in section 3 together with the conformal balanced condition on $M^{D}$ imply that the geometry of $M^{D}$ solves the heterotic supergravity Killing spinor equations (KSEs), i.e. $M^{D}$ is a (spacetime) supersymmetric background.
@@
In all cases, the gravitino KSE admits non-trivial solutions as a consequence of the restriction of the holonomy group of the $\widehat{\nabla}$ connection to one of those in (1). The same applies for the gaugino Killing spinor equation as a consequence of the condition (83) on the gauge coupling of the sigma model that arises from the requirement of the invariance of the sigma model action under the symmetries generated by the $\widehat{\nabla}$-covariantly constant forms associated with the holonomy group. Furthermore, the conformal balance condition always arises as one of the conditions of the dilatino KSE. So the target space of a conformally invariant sigma model whose couplings satisfy (81) and (83) is a supersymmetric background provided that the dilatino KSE does not impose additional conditions on the geometry of $M^{D}$ to that of conformal balance. To explore this statement requires a case by case investigation.
@@
First consider the CYT geometries for which the holonomy of $\widehat{\nabla}$ is contained in $SU(n)$. The dilatino KSE for such geometries implies that $M^{2n}$ is a conformally balanced complex manifold. As the integrability of the complex structure required by the dilatino KSE is already included in the definition of CYT structure, the target space of all conformally invariant sigma models with $\widehat{V}=0$ whose couplings satisfy the conditions required for the invariance of the sigma model action under the transformations generated by the $\widehat{\nabla}$-covariantly constant forms are (spacetime) supersymmetric. A similar conclusion holds for conformally balanced HKT and $\mathrm{Spin}(7)$ geometries.
@@
Next let us turn to the $G_{2}$ case. The dilatino KSE, apart from the conformal balance condition on $M^{7}$, it also imposes the condition that $H_{IJK}\varphi^{IJK}=0$. It is clear from the use of the Bianchi identity (54), the condition that the holonomy of $\widehat{\nabla}$ is included in $G_{2}$, the expression for ${\mathrm{d}}H$ in (12) and (72) that $H_{IJK}\varphi^{IJK}$ must be constant. The dilatino KSE implies that this constant vanish. This is an additional condition on the geometry which does not arise from either the conformal invariance of the sigma model or the invariance of the sigma model action under the symmetries generated by the $\widehat{\nabla}$-covariantly constant $G_{2}$ forms. Therefore, conformal invariance and the invariance of the sigma model actions under the symmetries generated by the $\widehat{\nabla}$-covariantly constant $G_{2}$ forms does not necessarily imply that the sigma model target space is supersymmetric. Though, we have not been able to provide an example to demonstrate that this obstruction to spacetime supersymmetry occurs.
@@ 4.3 Geometry of conformally invariant backgrounds with $\widehat{V}\not=0$
As $\widehat{V}$ is non-vanishing and $\widehat{\nabla}$-covariantly constant, the holonomy of $\widehat{\nabla}$ reduces further to a subgroup of the groups stated in (1). If the a priori holonomy group of $\widehat{\nabla}$ is one of the groups listed in (1), this subgroup will be specified as the invariance group of all $\widehat{\nabla}$-covariantly constant forms on $M^{D}$, including $\widehat{V}$. The pattern of holonomy reduction depends on the original a priori holonomy group of $\widehat{\nabla}$ and it will be outlined below in each case.
4.3.1 KT and CYT geometries and conformal invariance
Suppose that $M^{D}$ is a KT (or CYT) manifold that satisfies the conditions for conformal invariance with $\widehat{V}\not=0$. As it has already been mentioned the length of $\widehat{V}$ is constant, it is convenient in what follows to normalise $\widehat{V}$ to have length $1$,
@@ 4.3.1 KT and CYT geometries and conformal invariance
after possibly re-scaling it with an appropriate constant. As ^$\widehat{\nabla}\widehat{V}=0$ and ^$\widehat{\nabla}I=0$, one has that
^$\widehat{\nabla}\widehat{Z}=0~{},~{}~{}~{}\widehat{Z}^{J}=-I^{J}{}_{K}\widehat{V}^{K}~{}.$ (100)
The vector fields $\widehat{V}$ and $\widehat{Z}$ are orthogonal $G(\widehat{V},\widehat{Z})=0$ and $\widehat{Z}$ has length $1$ as well. Furthermore, $\widehat{Z}$, as $\widehat{V}$, leaves invariant $G$, $H$ and $F$ – the latter up to a gauge transformation. Indeed $\widehat{Z}$ is Killing as ^$\widehat{\nabla}\widehat{Z}=0$ and
@@
where we have used $\iota_{\widehat{Z}}F=0$ that in turn follows from $\iota_{\widehat{V}}F=0$ and that $F$ is (1,1)-form with respect to $I$. We have also taken $\iota_{\widehat{Z}}\widetilde{R}=0$ as we have done for $\widehat{V}$ in (96). The invariance of $F$ follows from $\iota_{\widehat{Z}}F=0$ and the argument stated in (98). One consequence of ^$\widehat{\nabla}\widehat{V}=\widehat{\nabla}\widehat{Z}=0$ is that the holonomy of $\widehat{\nabla}$ reduces to a subgroup of $U(n-1)$ in the KT case and to a subgroup of $SU(n-1)$ in the CYT case.
To continue, let us assume that the $[\widehat{V},\widehat{Z}]$ does not vanish. Using ${\mathrm{d}}\widehat{V}=\iota_{\widehat{V}}H$ and similarly for $\widehat{Z}$, one has
$[\widehat{V},\widehat{Z}]=\iota_{\widehat{V}}\iota_{\widehat{Z}}H~{},$ (102)
where it has been used that $G(\widehat{V},\widehat{Z})=0$. Therefore, the commutator of the two vector fields is given by components of $H$. As a consequence of this, the Bianchi identity (54), the condition (96) on $\widetilde{R}$ for both $\widehat{V}$ and $\widehat{Z}$ and (12), the commutator $[\widehat{V},\widehat{Z}]$ of $\widehat{V}$ with $\widehat{Z}$ is also a $\widehat{\nabla}$-covariantly constant vector field,
^$\widehat{\nabla}[\widehat{V},\widehat{Z}]=0~{}.$ (103)
As a result $[\widehat{V},\widehat{Z}]$ is Killing. It can also be shown that $\iota_{[\widehat{V},\widehat{Z}]}F=0$. Indeed,
@@
where we have used that $\iota_{\widehat{V}}F=\iota_{\widehat{Z}}F=0$ as well as the Bianchi identity for $F$. Consequently, $[\widehat{V},\widehat{Z}]$ leaves invariant both $H$ and $F$ – the latter up to a gauge transformation. The arguments for these are similar to those stated in (96) and (101), respectively. Furthermore, $[\widehat{V},\widehat{Z}]$ is orthogonal to both $\widehat{V}$ and $\widehat{Z}$, i.e.
@@
The above two equations can be demonstrated using that ^$\widehat{\nabla}G=\widehat{\nabla}I=0$ and (102) as well as the definition of $\widehat{Z}$. One consequence of the above relations is that after normalizing $[\widehat{V},\widehat{Z}]$ to have length $1$, as it was done for $\widehat{V}$, one can repeat the same process as that we have demonstrated with $\widehat{V}$ but now with $[\widehat{V},\widehat{Z}]$ in its place in the directions orthogonal to $\widehat{V}$ and $\widehat{Z}$. This means to consider the vector field $-I([\widehat{V},\widehat{Z}])$, which is again $\widehat{\nabla}$ covariantly constant, and observe that leaves invariant [^14] We also require that the condition (96) on $\widetilde{R}$ is satisfied with $\widehat{V}$ replaced by $[\widehat{V},\widehat{Z}]$. This is justified as $\breve{R}$ satisfies the same condition at the order of the perturbation theory considered. $G$, $H$ and $F$. This can be continued until the algebra of $\widehat{\nabla}$-covariantly constant vector fields closes to a Lie algebra $\mathfrak{g}$ of dimension $2p$. In such a case, the holonomy of $\widehat{\nabla}$ will reduce to $U(n-p)$ for KT geometries and to $SU(n-p)$ for CYT geometries.
As the vector fields generated by the action of $\mathfrak{g}$ on $M^{2n}$ have no fixed points, because they are nowhere vanishing, the action of $\mathfrak{g}$ on $M^{2n}$ is free. This can be integrated to an effective action of the (unique) group $\mathcal{G}$ on $M^{2n}$ which is the universal cover of all groups with Lie algebra $\mathfrak{g}$. The action of $\mathcal{G}$ on $M^{2n}$ is almost free, i.e. the isotropy group of any fixed points is a discrete subgroup of $\mathcal{G}$. This geometry is reminiscent of that encountered in the investigation of supersymmetric heterotic backgrounds in GLP and as a result we shall adopt a similar language to describe it. Assuming that $\mathcal{G}$, or at least an appropriate quotient of $\mathcal{G}$ with a discrete subgroup, acts freely (on the space of principal orbits) of $M^{2n}$, the geometry of $M^{2n}$ can be organised as follows: Let $(\lambda^{\alpha};a=1,\dots,2p)$ be the 1-forms dual to the $\widehat{\nabla}$-covariantly constant vector fields $(\widehat{V}_{\alpha};a=1,\dots,2p)$ generated by the action of $\mathfrak{g}$ on $M^{2n}$. One can view $(\lambda^{\alpha};\alpha=1,\dots,2p)$ as a principal bundle connection, $\lambda^{\alpha}(\widehat{V}_{\beta})=\delta^{\alpha}_{\beta}$. As $\lambda^{\alpha}$ is no-where vanishing and $G^{-1}(\lambda^{\alpha},\lambda^{\beta})=\delta^{\alpha\beta}$, because we have normalised $\widehat{V}_{\alpha}$ to have length $1$, it can be used to construct an orthonormal (co-)frame on $M^{2n}$ as $(\lambda^{\alpha},\mathbf{e}^{i};\alpha=1,\dots,2p,i=1,\dots 2n-2p)$. Then, the metric $G$ and $H$ can be written as
@@
where $H_{\alpha\beta\gamma}$ are the structure constants of $\mathfrak{g}$, $CS(\lambda)$ is the Chern-Simons form of $\lambda$, and $G^{\perp}=\delta_{ij}\mathbf{e}^{i}\mathbf{e}^{j}$ and $H^{\perp}={1\over 3!}H_{ijk}\mathbf{e}^{i}\wedge\mathbf{e}^{j}\wedge\mathbf{e}^{k}$ are the components of the metric and torsion orthogonal to all directions spanned by the vector fields generated by $\mathfrak{g}$ on $M^{2n}$. The component $H_{\alpha\beta i}$ of the torsion vanishes because of the closure of the algebra $\mathfrak{g}$ of $\widehat{\nabla}$-covariantly constant vector fields. Note that the Chern-Simon form of $\lambda$ is defined as

### v2

@@ 3.3 Worldsheet symmetry and geometry
$^\widehat{\nabla}$-covariantly constant forms on the spacetime $M^{D}$ generate symmetries in sigma model actions HowePap , like for example that of (2). More specifically, if the sigma model target space $M^{D}$ admits a $\widehat{\nabla}$ covariantly constant $(\ell+1)$-form $L$, i.e.
$\widehat{\nabla}_{I}L_{J_{1}\dots J_{\ell+1}}=0~,$ (81)
@@
where $\varepsilon_{L}$ is the infinitesimal parameter whose Grassmannian parity is chosen such that the variation $\delta_{L}X$ is even. It turns out that the conditions imposed on $F$ in (57), (66), (72) and (79) all follow from the condition (83) that arises from the requirement of invariance of heterotic sigma model action under the transformations (82) generated by the fundamental forms of the holonomy groups $SU(n)$, $Sp(k)$, $G_{2}$ and $\mathrm{Spin}(7)$ of $\widehat{\nabla}$ connection. This justifies the restrictions we have put on the curvature $F$ of the gauge sector in the discussion of geometries that satisfy the scale invariance conditions.
The commutator, $\delta_{L},\delta_{P}]X$ of two such transformations on $X$ generated by the $(\ell+1)$-form $L$ and the $(p+1)$-form $P$ on $M$ is rather contrived to be described here, see HPS for a detailed exposition. Even in the $G_{2}$ and ${\mathrm{S}pin}(7)$ cases, where the symmetries are generated by the fundamental forms that exhibit special algebraic properties, the expressions are rather involved, especially in the case that $H\not=0$. For the $G_{2}$ and ${\mathrm{S}pin}(7)$ geometries, the algebra of symmetries closes as a W-algebra, i.e. the structure constants of the algebra depend on conserved currents. The commutator on $\Psi$ is
$\delta_{L},\delta_{P}]\Psi_{-}^{a}=-A_{K}{}^{a}{}_{b}[\delta_{L},\delta_{P}]X^{K}-F_{IJ}{}^{a}{}_{b}\delta_{L}X^{I}\delta_{P}X^{J}\Psi_{-}^{b}~.$ (84)
@@
The simplest cases to consider are those for which the sigma-model target space admits either a KT structure II ($\widehat{\nabla}I=0$), or an HKT structure $I_{r}$ ($\widehat{\nabla}I_{r}=0$), then the action (2) is invariant under (0,2) or (0,4) worldsheet supersymmetry transformations [^12] There is an extensive literature, see for example Curtrightfreedman ; LAGDF ; Curtrightzachos ; Howesierra ; JGCHMR ; Buscher ; Braaten ; Hullreview ; HoweGP1 , on the geometry of target spaces of 2-dimensional sigma models with $(p,q)$, $p,q\not=0$, supersymmetry following the early work of Zumino2 in four dimensions., respectively. The additional supersymmetry transformations Hullwitten are given by
@@
where the integrability condition of the complex structures $I_{r}$ has been used [^13] The vanishing of the Nijenhuis tensor of $I_{r}$ simplifies the right hand side of the commutator (87). together with the conditions (83) for $L=I_{r}$. The latter conditions imply that $F$ is a (1,1)-form with respect to each complex structure $I_{r}$. In the KT case, the second term in the right hand side of the commutator can be expressed as $(\varepsilon D_{+}\varepsilon^{\prime}-\varepsilon^{\prime}D_{+}\varepsilon)D_{+}X^{K}$, while in the HKT case it can be expressed as $\delta_{rs}(\varepsilon^{r}D_{+}\varepsilon^{\prime s}-\varepsilon^{\prime s}D_{+}\varepsilon^{r})D_{+}X^{K}-(\varepsilon^{r}D_{+}\varepsilon^{\prime s}+\varepsilon^{\prime s}D_{+}\varepsilon^{r})\epsilon_{rs}{}^{t}I_{t}{}^{K}{}_{L}D_{+}X^{L}$. In either case, the commutator (87) closes to spacetime translations and supersymmetry transformations as expected. The commutator of two transformations (85) on $\Psi$ closes on-shell to the standard $(0,q)$ supersymmetry algebra, for either $q=2$ or $q=4$.
@@
where the superfields $X$ and $\Psi$ satisfy the constraint (88). It can be shown that this action is invariant under all $(0,q)$ supersymmetry transformations provided that in addition the fibre metric $h$ is (1,1), i.e. Hermitian, with respect to the fibre complex structures $J_{r}$, i.e. $h_{a^{\prime}b^{\prime}}J_{r}{}^{a^{\prime}}{}_{a}J_{r}{}^{b^{\prime}}{}_{b}=h_{ab}$ for either $r=1$ or $r=1,2,3$.
@@ 4.1 Comparison of conditions for scale and conformal invariance
$\widehat{\nabla}^{J}F_{JI}{}^{a}{}_{b}+\widehat{\theta}^{J}F_{JI}{}^{a}{}_{b}=0~,$ (91)
provided that (12) holds and the connection $\widetilde{\nabla}$ and $\mathcal{D}$} satisfy instanton-like conditions. A comparison of the above conditions for scale invariance with those for conformal invariance, which are the vanishing conditions for the beta functions of $G$, $B$ and $A$, in (21), reveals that these special geometries solve the latter provided that the 1-form
@@
$\widehat{\nabla}_{K}\widehat{V}_{L}=0~,~~~\widehat{V}^{K}F_{KL}{}^{a}{}_{b}=0\,$, (93)
where $\Phi$ is the dilaton. Therefore, $\widehat{V}$ is $^\widehat{\nabla}$-covariantly constant and the directions of $F$ along $\widehat{V}$ vanish. The former condition implies that
@@
Alternatively, suppose that $\widehat{V}\not=0$. As $\widehat{V}$ is non-vanishing and $^\widehat{\nabla}$-covariantly constant, it is non-vanishing everywhere on $M^{D}$. Moreover its length is constant as $\widehat{\nabla}$ is metric. The integrability condition of $\widehat{\nabla}\widehat{V}=0$ reveals that
@@
and the holonomy of the connection $\widehat{\nabla}$ reduces further to a subgroup of the original holonomy group stated in (1) whose geometry was summarised in sections 3.1 and 3.2. The pattern of these holonomy reductions depends on the original holonomy group and they will be explored below together with some aspects of the geometry of $M^{D}$. Here, we shall proceed to show that the couplings of the sigma model action are invariant under the diffeomorphisms generated by $\widehat{V}$.
@@
4.2 Conformal invariance with $\widehat{V}=0$ and spacetime supersymmetry
@@
We have demonstrated that the special geometries, investigated in section 3, that satisfy the conformal invariance conditions with $\widehat{V}=0$ are conformally balanced, i.e. $\widehat{\theta}=-2{\mathrm{d}}\Phi$. The conformal balance condition has arisen before in the context of (spacetime) supersymmetric heterotic backgrounds. Here, we shall address the question whether the invariance conditions (81) and (83) on the sigma model couplings, we have found in section 3, together with the conformal balanced condition on $M^{D}$ imply that the geometry of $M^{D}$ solves the heterotic supergravity Killing spinor equations (KSEs), i.e. $M^{D}$ is a (spacetime) supersymmetric background.
@@
In all cases, the gravitino KSE admits non-trivial solutions as a consequence of the restriction of the holonomy group of the $\widehat{\nabla}$ connection to be one of those in (1). The same applies for the gaugino Killing spinor equation. This is a consequence of the condition (83) on the gauge coupling of the sigma model that arises from the invariance of the sigma model action under the symmetries generated by the fundamental forms of the holonomy group. Furthermore, the conformal balance condition always arises as one of the conditions of the dilatino KSE. So the target space of a conformally invariant sigma model whose couplings satisfy (81) and (83) is a supersymmetric background provided that the dilatino KSE does not impose additional conditions on the geometry of $M^{D}$ to that of conformal balance. To explore this statement requires a case by case investigation.
@@
First consider the CYT geometries for which the holonomy of $\widehat{\nabla}$ is contained in $SU(n)$. The dilatino KSE for such geometries implies that $M^{2n}$ is a conformally balanced complex manifold. As the integrability of the complex structure required by the dilatino KSE is already included in the definition of CYT structure, the target space of all conformally invariant sigma models with $\widehat{V}=0$ whose couplings satisfy the conditions required for the invariance of the sigma model action under the transformations generated by the $^\widehat{\nabla}$-covariantly constant forms are (spacetime) supersymmetric. A similar conclusion holds for conformally balanced HKT and $\mathrm{Spin}(7)$ geometries.
@@
Next let us turn to the $G_{2}$ case. The dilatino KSE, apart from the conformal balance condition on $M^{7}$, it also imposes the condition that $H_{IJK}\varphi^{IJK}=0$. It is clear from the Bianchi identity (54), the condition that the holonomy of $\widehat{\nabla}$ is included in $G_{2}$, the expression for ${\mathrm{d}}H$ in (12) and (72) that $H_{IJK}\varphi^{IJK}$ must be constant. The dilatino KSE implies that this constant vanish. This is an additional condition on the geometry which does not arise from either the conformal invariance of the sigma model or the invariance of the sigma model action under the symmetries generated by the $^\widehat{\nabla}$-covariantly constant $G_{2}$ forms. Therefore, conformal invariance and the invariance of the sigma model actions under the symmetries generated by the $^\widehat{\nabla}$-covariantly constant $G_{2}$ forms does not necessarily imply that the sigma model target space is supersymmetric. This phenomenon occurs and an example will be provided in section 4.4.3 below.
@@ 4.3 Geometry of conformally invariant backgrounds with $\widehat{V}\not=0$
As $\widehat{V}$ is non-vanishing and $^\widehat{\nabla}$-covariantly constant vector field, the holonomy of $\widehat{\nabla}$ reduces further to a subgroup of the a priori holonomy groups stated in (1). This subgroup is specified as the isotropy group of all $^\widehat{\nabla}$-covariantly constant forms on $M^{D}$, including $\widehat{V}$. The pattern of holonomy reduction depends on the original a priori holonomy group of $\widehat{\nabla}$ and it will be outlined below on a case by case basis.
4.3.1 KT geometries with $^\widehat{\nabla}$-covariantly constants vector fields.
Suppose that $M^{D}$ is a KT (or CYT) manifold that satisfies the conditions for conformal invariance with $\widehat{V}\not=0$. As the length of $\widehat{V}$ is constant, because $\widehat{V}$ is $^\widehat{\nabla}$-covariantly constant, it is convenient in what follows to normalise $\widehat{V}$ to have length $1$,
@@
after possibly an appropriate constant re-scaling. As $\widehat{\nabla}\widehat{V}=0$ and $\widehat{\nabla}I=0$, one has that
$\widehat{\nabla}\widehat{Z}=0~,~~~\widehat{Z}^{J}=-I^{J}{}_{K}\widehat{V}^{K}~.$ (100)
The vector fields $\widehat{V}$ and $\widehat{Z}$ are orthogonal $G(\widehat{V},\widehat{Z})=0$ and $\widehat{Z}$ has length $1$ as well. Furthermore, $\widehat{Z}$, as $\widehat{V}$, leaves invariant $G$, $H$ and $F$ – the latter up to a gauge transformation. Indeed, $\widehat{Z}$ is Killing as $\widehat{\nabla}\widehat{Z}=0$ and leave $H$ invariant as
@@
where we have used $\iota_{\widehat{Z}}F=0$. In turn, this follows from $\iota_{\widehat{V}}F=0$ and the property of $F$ to be an (1,1)-form with respect to $I$. We have also used $\iota_{\widehat{Z}}\widetilde{R}=0$ as we have done for $\widehat{V}$ in (96). The invariance of $F$ follows from $\iota_{\widehat{Z}}F=0$ and the argument stated in (98). One consequence of $\widehat{\nabla}\widehat{V}=\widehat{\nabla}\widehat{Z}=0$ is that the holonomy of $\widehat{\nabla}$ reduces from $U(n)$ to the subgroup U(n1)U(n-1) in the KT case and from $SU(n)$ to the subgroup SU(n1)SU(n-1) in the CYT case.
To continue, let us assume that the $\widehat{V},\widehat{Z}$] does not vanish. Using ${\mathrm{d}}\widehat{V}=\iota_{\widehat{V}}H$ and similarly for $\widehat{Z}$, one has
$\widehat{V},\widehat{Z}]=\iota_{\widehat{V}}\iota_{\widehat{Z}}H~,$ (102)
where it has been used that $G(\widehat{V},\widehat{Z})=0$. Therefore, the commutator of the two vector fields is given by components of $H$. As a consequence of this, the Bianchi identity (54), the condition (96) on $\widetilde{R}$ for both $\widehat{V}$ and $\widehat{Z}$ and (12), the commutator $\widehat{V},\widehat{Z}$] of $\widehat{V}$ with $\widehat{Z}$ is also a $^\widehat{\nabla}$-covariantly constant vector field,
$\widehat{\nabla}[\widehat{V},\widehat{Z}]=0~.$ (103)
As a result $\widehat{V},\widehat{Z}$] is Killing. It can also be shown that $\iota_{[\widehat{V},\widehat{Z}]}F=0$. Indeed,
@@
where we have used that $\iota_{\widehat{V}}F=\iota_{\widehat{Z}}F=0$ as well as the Bianchi identity for $F$. Consequently, $\widehat{V},\widehat{Z}$] leaves invariant both $H$ and $F$ – the latter up to a gauge transformation. The arguments for these are similar to those stated in (96) and (101), respectively. Furthermore, $\widehat{V},\widehat{Z}$] is orthogonal to both $\widehat{V}$ and $\widehat{Z}$, i.e.
@@
[…]
As the vector fields generated by the action of $\mathfrak{g}$ on $M^{2n}$ have no fixed points, because they are nowhere vanishing, the action of $\mathfrak{g}$ on $M^{2n}$ is free. This can be integrated to an effective action of the (unique) group $\mathcal{G}$ on $M^{2n}$ which is the universal cover of all groups with Lie algebra $\mathfrak{g}$. The action of $\mathcal{G}$ on $M^{2n}$ is almost free, i.e. the isotropy group of any fixed points is a discrete subgroup of $\mathcal{G}$. This geometry is reminiscent of that encountered in the investigation of supersymmetric heterotic backgrounds in GLP and as a result we shall adopt a similar language to describe it. Assuming that $\mathcal{G}$, or at least an appropriate quotient of $\mathcal{G}$ with a discrete subgroup, acts freely (on the space of principal orbits) of $M^{2n}$, the geometry of $M^{2n}$ can be organised as follows: Let $(\lambda^{\alpha};a=1,\dots,2p)$ be the 1-forms dual to the $^\widehat{\nabla}$-covariantly constant vector fields $(\widehat{V}_{\alpha};a=1,\dots,2p)$ generated by the action of $\mathfrak{g}$ on $M^{2n}$. One can view $(\lambda^{\alpha};\alpha=1,\dots,2p)$ as a principal bundle connection, $\lambda^{\alpha}(\widehat{V}_{\beta})=\delta^{\alpha}_{\beta}$. As $\lambda^{\alpha}$ is no-where vanishing and $G^{-1}(\lambda^{\alpha},\lambda^{\beta})=\delta^{\alpha\beta}$, because we have normalised $\widehat{V}_{\alpha}$ to have length $1$, it can be used to construct an orthonormal (co-)frame on $M^{2n}$ as $(\lambda^{\alpha},\mathbf{e}^{i};\alpha=1,\dots,2p,i=1,\dots 2n-2p)$. Then, the metric $G$ and $H$ can be written as
@@
where $H_{\alpha\beta\gamma}$ are the structure constants of $\mathfrak{g}$, $CS(\lambda)$ is the Chern-Simons form of $\lambda$, and $G^{\perp}=\delta_{ij}\mathbf{e}^{i}\mathbf{e}^{j}$ and $H^{\perp}={1\over 3!}H_{ijk}\mathbf{e}^{i}\wedge\mathbf{e}^{j}\wedge\mathbf{e}^{k}$ are the components of the metric and torsion orthogonal to all directions spanned by the vector fields generated by $\mathfrak{g}$ on $M^{2n}$. The component $H_{\alpha\beta i}$ of the torsion vanishes because of the closure of the algebra $\mathfrak{g}$ of $^\widehat{\nabla}$-covariantly constant vector fields. Note that the Chern-Simon form of $\lambda$ is defined as